Given the following semiannual-pay bond:Corporate Bond, non-callable, Matures in 10 years, 5% Annual Coupon, YTM 7%, purchases $1,000 in Face Value. What is the present value of the coupon portion of cash flows? (1). 257.82 (2). 355.31 (3) 397.40 (4) 426.43
Answer #1

Coupon = (5% of 1000) / 2 = 25

Number of periods = 10 * 2= 20

Rate = 7% / 2 = 3.5

Present value = Coupon * [1 - 1 / (1 + r)n]/ r

Present value = 25 * [1 - 1 / (1 + 0.035)20]/ 0.035

Present value = 25 * [1 - 0.502566] / 0.035

Present value = 25 * 14.212403

Present value = $355.31